# Optimizing Generative Ranking Relevance via Reinforcement Learning in Xiaohongshu Search

### v1

@@ Abstract.
Ranking relevance is a fundamental task in search engines, aiming to identify the items most relevant to a given user query. Traditional relevance models typically produce scalar scores or directly predict relevance labels, limiting both interpretability and the modeling of complex relevance signals. Inspired by recent advances in Chain-of-Thought (CoT) reasoning for complex tasks, we investigate whether explicit reasoning can enhance both interpretability and performance in relevance modeling. However, existing reasoning-based Generative Relevance Models (GRMs) primarily rely on supervised fine-tuning on large amounts of human-annotated or synthetic CoT data, which often leads to limited generalization. Moreover, domain-agnostic, free-form reasoning tends to be overly generic and insufficiently grounded, limiting its potential to handle the diverse and ambiguous cases prevalent in open-domain search. In this work, we formulate relevance modeling in Xiaohongshu search as a reasoning task and introduce a Reinforcement Learning (RL)–based training framework to enhance the grounded reasoning capabilities of GRMs. Specifically, we incorporate practical business-specific relevance criteria into the multi-step reasoning prompt design and propose Stepwise Advantage Masking (SAM), a lightweight process-supervision strategy which facilitates effective learning of these criteria through improved credit assignment. To enable industrial deployment, we further distill the large-scale RL-tuned model to a lightweight version suitable for real-world search systems. Extensive experiments on industrial datasets, along with online A/B tests, demonstrate the effectiveness of our approach.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Similarity measuresccs: Information systems Language models
@@ 1. Introduction
Search engines (croft2010search) such as Google and Xiaohongshu [^1] https://www.xiaohongshu.com/explore have become indispensable gateways to online information, serving hundreds of millions of user queries every day. A key enabler of their effectiveness is the accurate assessment of ranking relevance (YahooSearch), which fundamentally underpins improvements in user experience and satisfaction. Traditional relevance modeling approaches predominantly adopt scalar discriminative models (bertrank; yates-etal-2021-pretrained; BaiduSearch), producing a single probability that indicates whether a query–item pair is relevant. Despite their success, these models operate as “black boxes”, offering limited interpretability and making it difficult to diagnose relevance errors. Moreover, their limited knowledge capacity restricts their ability to understand the complex and nuanced semantics underlying relevance.
@@ 2.1. Generative Relevance Models
The rapid advancement of LLMs has shifted relevance modeling from traditional discriminative classification (bertrank; yates-etal-2021-pretrained; rankt5) to generative paradigms (i.e., GRMs) (zhuang-etal-2024-beyond; sun-etal-2023-chatgpt). GRMs exploit the text generation capabilities of LLMs to assess query–item relevance in a more flexible and interpretable manner. Early GRMs follow a pointwise scoring strategy, deriving relevance from the likelihood of specific token outputs (e.g., “Yes” or “No”) (liang2023holistic; zhuang-etal-2024-beyond). Recent progress introduces ranking-oriented prompting strategies that support pairwise (qin-etal-2024-large) and listwise (sun-etal-2023-chatgpt) comparisons, benefiting from LLMs’ instruction-following abilities. To further enhance relevance assessment, CoT prompting has been incorporated into GRMs (niu2024judgerankleveraginglargelanguage; weller2025rank1testtimecomputereranking), enabling multi-step reasoning akin to human judgment. However, existing reasoning-based approaches either depend on SFT with costly collected reasoning paths (Ecommerce2; Ecommerce1; weller2025rank1testtimecomputereranking), which may generalize poorly to long-tail search queries, or adopt outcome-based RL (zhuang2025rankr1enhancingreasoningllmbased) without fine-grained supervision of intermediate reasoning steps, leading to suboptimal reasoning chains. In contrast, we introduce a process-supervised RL framework that performs step-level credit assignment, yielding more grounded reasoning trajectories and significantly improving robustness in ambiguous search scenarios.
@@ 3. Preliminaries
Relevance modeling in Xiaohongshu search is defined as a multi-class classification task, whose objective is to assign a discrete relevance label $y\in\{-1,0,1,2,3\}$ to each query–note pair $(q,n)$. Each label corresponds to a distinct level of semantic relevance, ranging from strongly irrelevant (1-1), irrelevant (0), weakly relevant ($1$), partially relevant ($2$), to perfectly relevant ($3$). The annotation scheme is derived from human judgments made by trained evaluators, following detailed guidelines that account for search intent alignment, topical consistency, expected user engagement, and other relevance-related dimensions.
@@ 4. Methodology
This section outlines our RL–based training paradigm. Section 4.1 reformulates relevance modeling from a generative perspective, enabling the model to jointly produce reasoning traces and final assessments. Section 3 introduces our criteria-augmented prompt design, while Section 4.3 presents a distillation-driven cold-start initialization that stabilizes subsequent RL training. Finally, Section 4.4 details our process-supervised RL framework and its Stepwise Advantage Masking (SAM) mechanism, which further improves reasoning fidelity and quality.
@@ 4.2. Criteria-augmented Prompt Design
While recent advances in RL with verified rewards have demonstrated strong capabilities in mathematical reasoning (r1) and code generation (dou-etal-2024-stepcoder), extending this paradigm to search relevance assessment remains fundamentally challenging. Mathematical and coding tasks are inherently objective, governed by universal axioms and formal logic. Since LLMs already acquire the foundational knowledge for such tasks during pretraining, RL can further enhance performance by promoting unconstrained reasoning grounded in this prior (yue2025does). In contrast, relevance assessment is intrinsically subjective and domain-specific. The relevance criteria are tightly coupled with the specific business context and user ecosystem of platforms like Xiaohongshu, especially in ambiguous search cases. Without specific guidance, a model cannot reliably infer such nuanced criteria through unguided exploration, often leading to suboptimal performance. To bridge this gap, we leverage Xiaohongshu’s relevance criteria, a structured body of operational knowledge developed over years of search optimization. These criteria include general relevance principles defined by human experts, as well as specialized rules for handling edge cases, which together serve as pseudo-axioms to guide the model’s reasoning process. Building on these, we design a structured prompt (see Figure 3) that internally decomposes the relevance assessment into three interpretable reasoning steps:
@@ 5.1.1. Benchmarks and Metrics
Our experiments are conducted on two benchmark datasets widely used in real-world search deployments: RANDOM and LONGTAIL. Both datasets consist of 15,000 manually annotated query–note pairs, but differ in their query sampling distributions, enabling us to evaluate model robustness under both common and long-tail user behaviors. The statistics of these datasets are summarized in Table 1. We provide a detailed description of each dataset below:
  • RANDOM: The queries are randomly sampled from a large-scale online search system, and for each query, notes are selected from candidate note lists at various ranking positions. This design ensures that the dataset closely reflects the naturally occurring data distribution observed in practical deployments. Each query–note pair is manually annotated and has undergone multiple rounds of quality inspection.
@@ 5.1.2. Data Preparation
To construct the training dataset for our framework, we adopt the same query and note sampling strategy as used in the RANDOM benchmark. In total, we sampled 50K queries, each paired with 20 notes. 25 professionally trained annotators labeled 1M query–note pairs over three months. The overall label distribution is consistent with that of the RANDOM dataset. To obtain the initial CoT reasoning data, we employ DeepSeek-R1 (r1) to generate reasoning traces using the proposed criteria-augmented reasoning prompt. Given that the raw reasoning outputs may be unreliable, we apply rejection sampling to retain only those traces whose final predictions are consistent with human annotations. Since the rejection sampling step inevitably alters the original label distribution, we slightly over-sample or down-sample certain instances to restore it to match the original dataset. This process yields 500K validated reasoning trajectories. To further enhance data quality in RL training, we introduce an offline difficulty estimation strategy based on avg@k. For each query–note instance, we compute avg@k as the average prediction accuracy obtained by sampling the initial policy reasoning path $k$ times:
@@ 5.2.3. Model dynamics between SFT and RL
As shown in Table 3, ProcessRL-Reasoning demonstrates superior F1-scores across most categories compared to SFT-label with the same dataset size of 200k. In the 2-class setting, the improvement is particularly notable in the precision for the “0*” category (i.e., the aggregation of labels -1/0) and the recall for the “1*” category (i.e., the aggregation of labels 1/2/3), where the precision rate increases from 82.67 to 90.31, and the recall rate improves from 86.70 to 93.45. This difference highlights that ProcessRL-Reasoning is more cautious in identifying irrelevant examples and more proactive in recognizing potentially relevant ones, significantly addressing the under-recall issue observed in SFT-label when handling ambitious positive examples. From the perspective of the 5-class setting, the recall for the “1*” category improves through a general enhancement across labels 1/2/3, with only a slight sacrifice in precision for labels 2 and 3. Furthermore, ProcessRL-Reasoning shows significant improvements in both precision and recall for the minority classes (labels -1 and 1), demonstrating its strong ability to handle long-tail cases. This improvement in both precision and recall indicates that our method refines the decision boundary by optimizing the reasoning process itself, rather than simply learning input-output mappings. These results suggest that incorporating grounded reasoning—where the model relies on domain-specific relevance criteria—enables the model to capture subtle distinctions between fine-grained categories, leading to improved classification performance.
@@ 5.3.1. Online Deployment
To reduce inference latency in online deployment, we employ a knowledge distillation strategy (44873). Specifically, we first use our RL-tuned model (i.e., the teacher) to annotate millions of samples, retaining only the final-step scores as labels. We then distill these capabilities into a lightweight 0.1B-parameter BERT-based model (i.e., the student), which enables efficient online inference. In production, the deployed student model consistently achieves a P95 latency of around 20ms and maintains an almost 100% response success rate. Its predicted relevance scores are forwarded to the next ranking module, which determines the final display order of notes in the Xiaohongshu app. To further enhance user experience, we adopt a binary classification setting, tagging each note as either relevant or irrelevant. Relevant notes are prioritized in the ranking results, while irrelevant ones are relegated to lower positions.
@@ 5.3.2. Online Performance
To evaluate the student model, we deployed it on the Xiaohongshu A/B testing platform. We randomly sampled 5% of the online traffic as the test group and applied our proposed approach to this group. Another 5% of the traffic was used as the baseline group, which continued to rely on the previous BERT-based model. For a fair comparison, we continuously monitored the experimental results on the platform. To mitigate the impact of traffic fluctuations, we set a minimum testing period of seven days. Notably, we adopt the following two business metrics to evaluate online user experience:
@@
According to the results in Table 4, our approach achieves a 0.72% improvement in the CES metric, reflecting increased user engagement and an enhanced user experience. In addition, the DCG 0/1 metric shows a cumulative reduction of 0.36%, indicating fewer irrelevant results and improved ranking quality. Overall, these findings demonstrate that our approach significantly enhances both user experience and retrieval relevance.

### v2

@@ Abstract.
Ranking relevance is a fundamental task in search engines, aiming to identify the items most relevant to a given user query. Traditional relevance models typically produce scalar scores or directly predict relevance labels, limiting both interpretability and the modeling of complex relevance signals. Inspired by recent advances in Chain-of-Thought (CoT) reasoning for complex tasks, we investigate whether explicit reasoning can enhance both interpretability and performance in relevance modeling. However, existing reasoning-based Generative Relevance Models (GRMs) primarily rely on supervised fine-tuning on large amounts of human-annotated or synthetic CoT data, which often leads to limited generalization. Moreover, domain-agnostic, free-form reasoning tends to be overly generic and insufficiently grounded, limiting its potential to handle the diverse and ambiguous cases prevalent in open-domain search. In this work, we formulate relevance modeling in Xiaohongshu search as a reasoning task and introduce a Reinforcement Learning (RL)-based training framework to enhance the grounded reasoning capabilities of GRMs. Specifically, we incorporate practical business-specific relevance criteria into the multi-step reasoning prompt design and propose Stepwise Advantage Masking (SAM), a lightweight process-supervision strategy which facilitates effective learning of these criteria through improved credit assignment. To enable industrial deployment, we further distill the large-scale RL-tuned model to a lightweight version suitable for real-world search systems. Extensive offline evaluations and online A/B tests demonstrate that our approach consistently delivers significant improvements across key relevance and business metrics, validating its effectiveness, robustness, and practicality for large-scale industrial search systems.
@@
copyright: acmlicensedjournalyear: 2026copyright: ccconference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 09–13, 2026; Jeju Island, Republic of Koreabooktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’26), August 09–13, 2026, Jeju Island, Republic of Koreadoi: 10.1145/3770854.3783917isbn: 979-8-4007-2258-5/2026/08ccs: Information systems Similarity measuresccs: Information systems Language models
@@ 1. Introduction
Search engines (croft2010search) such as Google and Xiaohongshu [^1] https://www.xiaohongshu.com/explore have become indispensable gateways to online information, serving hundreds of millions of user queries every day. A key enabler of their effectiveness is the accurate assessment of ranking relevance (YahooSearch), which fundamentally underpins improvements in user experience and satisfaction. Traditional relevance modeling approaches predominantly adopt scalar discriminative models (bertrank; yates-etal-2021-pretrained; BaiduSearch), producing a single probability that indicates whether a query-item pair is relevant. Despite their success, these models operate as “black boxes”, offering limited interpretability and making it difficult to diagnose relevance errors. Moreover, their limited knowledge capacity restricts their ability to understand the complex and nuanced semantics underlying relevance.
@@ 2.1. Generative Relevance Models
The rapid advancement of LLMs has shifted relevance modeling from traditional discriminative classification (bertrank; yates-etal-2021-pretrained; rankt5) to generative paradigms (i.e., GRMs) (zhuang-etal-2024-beyond; sun-etal-2023-chatgpt). GRMs exploit the text generation capabilities of LLMs to assess query-item relevance in a more flexible and interpretable manner. Early GRMs follow a pointwise scoring strategy, deriving relevance from the likelihood of specific token outputs (e.g., “Yes” or “No”) (liang2023holistic; zhuang-etal-2024-beyond). Recent progress introduces ranking-oriented prompting strategies that support pairwise (qin-etal-2024-large) and listwise (sun-etal-2023-chatgpt) comparisons, benefiting from LLMs’ instruction-following abilities. To further enhance relevance assessment, CoT prompting has been incorporated into GRMs (niu2024judgerankleveraginglargelanguage; weller2025rank1testtimecomputereranking), enabling multi-step reasoning akin to human judgment. However, existing reasoning-based approaches either depend on SFT with costly collected reasoning paths (Ecommerce2; Ecommerce1; weller2025rank1testtimecomputereranking), which may generalize poorly to long-tail queries, or adopt outcome-based RL (zhuang2025rankr1enhancingreasoningllmbased; zhang-etal-2025-rearank) without fine-grained supervision of intermediate reasoning steps, leading to suboptimal reasoning chains. In contrast, we introduce a process-supervised RL framework that performs step-level credit assignment, yielding more grounded reasoning trajectories and significantly improving robustness in ambiguous search scenarios.
@@ 3. Preliminaries
Relevance modeling in Xiaohongshu search is defined as a multi-class classification task, whose objective is to assign a discrete relevance label $y\in\{-1,0,1,2,3\}$ to each query-note pair $(q,n)$. Each label corresponds to a distinct level of semantic relevance, ranging from strongly irrelevant (1-1), irrelevant (0), weakly relevant ($1$), partially relevant ($2$), to perfectly relevant ($3$). The annotation scheme is derived from human judgments made by trained evaluators, following detailed guidelines that account for search intent alignment, topical consistency, expected user engagement, and other relevance-related dimensions.
@@ 4. Methodology
This section outlines our RL-based training paradigm. Section 4.1 reformulates relevance modeling from a generative perspective, enabling the model to jointly produce reasoning traces and final assessments. Section 3 introduces our criteria-augmented prompt design, while Section 4.3 presents a distillation-driven cold-start initialization that stabilizes subsequent RL training. Finally, Section 4.4 details our process-supervised RL framework and its Stepwise Advantage Masking (SAM) mechanism, which further improves reasoning fidelity and quality.
@@ 4.2. Criteria-augmented Prompt Design
While recent advances in RL with verified rewards have demonstrated strong capabilities in mathematical reasoning (r1), extending this paradigm to search relevance assessment remains fundamentally challenging. Mathematical and coding tasks are inherently objective, governed by universal axioms and formal logic. Since LLMs already acquire the foundational knowledge for such tasks during pretraining, RL can further enhance performance by promoting unconstrained reasoning grounded in this prior (yue2025does). In contrast, relevance assessment is intrinsically subjective and domain-specific. The relevance criteria are tightly coupled with the specific business context and user ecosystem of platforms like Xiaohongshu, especially in ambiguous search cases. Without specific guidance, a model cannot reliably infer such nuanced criteria through unguided exploration, often leading to suboptimal performance. To bridge this gap, we leverage Xiaohongshu’s relevance criteria, a structured body of operational knowledge developed over years of search optimization. These criteria include general relevance principles defined by human experts, as well as specialized rules for handling edge cases, which together serve as pseudo-axioms to guide the model’s reasoning process. Building on these, we design a structured prompt (see Figure 3) that internally decomposes the relevance assessment into three interpretable reasoning steps:
@@ 5.1.1. Benchmarks and Metrics
Our experiments are conducted on two benchmark datasets widely used in real-world search deployments: RANDOM and LONGTAIL. Both datasets consist of 15,000 manually annotated query-note pairs, but differ in their query sampling distributions, enabling us to evaluate model robustness under both common and long-tail user behaviors. The statistics of these datasets are summarized in Table 1. We provide a detailed description of each dataset below:
  • RANDOM: The queries are randomly sampled from a large-scale online search system, and for each query, notes are selected from candidate note lists at various ranking positions. This design ensures that the dataset closely reflects the naturally occurring data distribution observed in practical deployments. Each query-note pair is manually annotated and has undergone multiple rounds of quality inspection.
@@ 5.1.2. Data Preparation
To construct the training dataset for our framework, we adopt the same query and note sampling strategy as used in the RANDOM benchmark. In total, we sampled 50K queries, each paired with 20 notes. 25 professionally trained annotators labeled 1M query-note pairs over three months. The overall label distribution is consistent with that of the RANDOM dataset. To obtain the initial CoT reasoning data, we employ DeepSeek-R1 (r1) to generate reasoning traces using the proposed criteria-augmented reasoning prompt. Given that the raw reasoning outputs may be unreliable, we apply rejection sampling to retain only those traces whose final predictions are consistent with human annotations. Since the rejection sampling step inevitably alters the original label distribution, we slightly over-sample or down-sample certain instances to restore it to match the original dataset. This process yields 500K validated reasoning trajectories. To further enhance data quality in RL training, we introduce an offline difficulty estimation strategy based on avg@k. For each query–note instance, we compute avg@k as the average prediction accuracy obtained by sampling the initial policy reasoning path $k$ times:
@@ 5.2.3. Model dynamics between SFT and RL
As shown in Table 3, ProcessRL-Reasoning demonstrates superior F1-scores across most categories compared to SFT-label with the same dataset size of 200k. In the 2-class setting, the improvement is particularly notable in the precision for the “0*” category (i.e., the aggregation of labels -1/0) and the recall for the “1*” category (i.e., the aggregation of labels 1/2/3), where the precision rate increases from 82.67 to 90.31, and the recall rate improves from 86.70 to 93.45. This difference highlights that ProcessRL-Reasoning is more cautious in identifying irrelevant examples and more proactive in recognizing potentially relevant ones, significantly addressing the under-recall issue observed in SFT-label when handling ambiguous positive examples. From the perspective of the 5-class setting, the recall for the “1*” category improves through a general enhancement across labels 1/2/3, with only a slight sacrifice in precision for labels 2 and 3. Furthermore, ProcessRL-Reasoning shows significant improvements in both precision and recall for the minority classes (labels -1 and 1), demonstrating its strong ability to handle long-tail cases. This improvement in both precision and recall indicates that our method refines the decision boundary by optimizing the reasoning process itself, rather than simply learning input-output mappings. These results suggest that incorporating grounded reasoning—where the model relies on domain-specific relevance criteria—enables the model to capture subtle distinctions between fine-grained categories, leading to improved classification performance.
@@ 5.3.1. Online Deployment
To reduce inference latency in online deployment, we adopt a knowledge distillation strategy (44873). Specifically, we use our RL-tuned model as a teacher to annotate millions of query–note pairs, and retain only the final-step relevance scores as supervision signals. These signals are then distilled into a lightweight 0.1B-parameter BERT-based discriminative 5-class classifier, which serves as the student model for online inference. In production, the deployed student model achieves a P95 latency of approximately 20ms with an almost 100% response success rate. To preliminarily assess the effectiveness of the distilled student model, we conduct an offline evaluation on the RANDOM test set. The student model attains a 2-ACC of 90.65, a 5-ACC of 79.22, a Macro F1 of 68.19, and a weighted F1 of 78.57. These results consistently outperform the SFT-Label baseline, while still exhibiting a noticeable gap compared to the teacher model. This gap indicates the significant room for further improvement through more advanced distillation strategies.
@@ 5.3.2. Online Performance
To evaluate the student model online, we deployed it on the Xiaohongshu A/B testing platform. We randomly sampled 5% of the online traffic as the test group and applied our proposed approach to this group. Another 5% of the traffic was used as the baseline group, which continued to rely on the previous BERT-based model. For a fair comparison, we continuously monitored the experimental results on the platform. To mitigate the impact of traffic fluctuations, we set a minimum testing period of seven days. Notably, we adopt the following two business metrics to evaluate online user experience:
@@
According to the results in Table 4, our approach achieves a 0.72% improvement in the CES metric, reflecting increased user engagement and an enhanced user experience. In addition, the DCG 0/1 metric shows a cumulative reduction of 0.36%, indicating fewer irrelevant results and improved ranking quality. Overall, these findings demonstrate that our approach significantly enhances both user experience and ranking relevance.